OpenStudy (anonymous):

1/2a^2 - 1/4a For a=4

OpenStudy (lgbasallote):

substitute a with 4....where are you confused tehre?

OpenStudy (anonymous):

1/2(4)^2 - 1/4(4) 1/2 (16) - 1/4(4) 16/2- 4/4 ???

OpenStudy (lgbasallote):

that depends... is the question \[\frac 12 a^2 - \frac 14a\]or\[\frac 1{2a^2} - \frac 1{4a}\]

OpenStudy (anonymous):

its 1/2a2 - 1/4a

OpenStudy (lgbasallote):

?

OpenStudy (anonymous):

a^2

OpenStudy (lgbasallote):

a^2 = \(a^2\)

OpenStudy (anonymous):

I am working on polynomials. The question says to "Evaluate the expression" 1/2a^2 - 1/4a For a= 4

OpenStudy (lgbasallote):

and i am asking you how the question looks like...

OpenStudy (anonymous):

How do you get your equation to look like a real exponent on here?

OpenStudy (lgbasallote):

is the a in the denominator or not?

OpenStudy (lgbasallote):

is it \[\huge \frac 12 a^2 - \frac 14a\] or \[\huge \frac 1{2a^2} - \frac 1{4a}\]

OpenStudy (anonymous):

no, it is to the side of the of the fraction

OpenStudy (lgbasallote):

then you're right

OpenStudy (anonymous):

Got it.... thank you- but seriously. how did you type out the equation like that on the computer

OpenStudy (lgbasallote):

do you see that box that says \[\huge \boxed{\sum \; \rm{Equation}}\]

OpenStudy (lgbasallote):

below the comment box

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh.. cool.. thank you

OpenStudy (lgbasallote):

welcome